# A Roadmap for Software Testing in Open Collaborative Development Environments

### v1

@@ 5.3. Quality Assurance for Auto-generated Code
With the advancement of LLMs and their remarkable performance in code generation tasks, developers are increasingly relying on these models for various coding and debugging tasks. Recently, there have been reports of groundbreaking developments, such as Microsoft’s creation of an AI programmer named AutoDev (Microsoft, 2024), capable of mastering full-stack skills, which can not only write code and debug, but also train models and even bid for projects on the largest freelancing platform, Upwork. As a result, the proliferation of AI-generated code inevitably finds its way into the open-source community, presenting significant challenges for ensuring the overall quality of the open-source ecosystem.
The integration of AI-generated code into open-source projects raises concerns regarding code reliability, maintainability, and security. While AI models demonstrate impressive capabilities in code generation, there remain uncertainties about the robustness and correctness of the generated code (Ouyang et al., 2023). Additionally, the lack of human oversight in the code generation process may lead to the introduction of vulnerabilities or inefficiencies that could compromise the integrity of software systems. Furthermore, the rapid adoption of AI-generated code in open-source projects exacerbates the challenge of maintaining code quality standards across diverse programming languages and development environments. The sheer volume of AI-generated contributions necessitates scalable quality assurance processes to ensure that the code meets established coding standards, adheres to best practices, and complies with project-specific requirements.
To address these challenges, it becomes imperative to develop comprehensive quality assurance frameworks tailored specifically for AI-generated code. Such frameworks should encompass automated code review mechanisms, static analysis tools, and robust testing methodologies to detect and rectify potential issues in the generated code. Moreover, fostering collaboration between AI developers, software engineers, and open-source maintainers is essential to facilitate the seamless integration of AI-generated contributions while upholding the quality and reliability of open-source software projects.
5.4. AI Models Ecosystem and Security
In addition to platforms like GitHub, which host a plethora of code contributions from programmers worldwide, the advancement of AI technology has elevated AI models to crucial components in software development. Many developers now rely on pre-trained AI models downloaded from repositories or fine-tune these models with domain-specific data to achieve specific outcomes. This trend has spurred the emergence of platforms such as Hugging Face (Face, 2024), which serves as repositories for AI models and facilitate their distribution and usage.
In such platforms, where AI models are distributed and utilized, various general security risks may arise. For example, data poisoning can inject manipulated data during model fine-tuning, potentially leading to biased outcomes (Shafahi et al., 2018), while backdoor attacks entail covertly implanting vulnerabilities into AI models, compromising their integrity (Liu et al., 2018). Ensuring the quality of public AI models hosted on these platforms is essential to foster trust and reliability among users and promote the widespread adoption of AI technology. This includes secure data handling practices to protect privacy, implementing mechanisms for model authentication and integrity verification, and deploying defense mechanisms against adversarial attacks, such as robust model testing and validation techniques. Moreover, fostering a culture of security awareness and promoting collaboration between security experts and AI practitioners can help identify and address potential vulnerabilities effectively.
5.5. Process Management and People Collaboration in AI Model-centric Paradigm
The emerging trend of utilizing AI models, downloaded, fine-tuned, and shared on platforms like Hugging Face, may necessitate additional forms of process management and collaboration among individuals to effectively support the development workflow.
In such a scenario, traditional process management methodologies may face challenges in adapting to the dynamic and decentralized nature of AI model-centric open collaboration. One potential challenge is maintaining version control and ensuring consistency across different iterations of AI models, especially when multiple contributors are involved in fine-tuning and uploading. Additionally, ensuring the quality and reliability of AI models may require new strategies for testing, validation, and integration into existing systems. Moreover, effective collaboration within the context of AI model-centric open collaboration may also face challenges. One potential challenge arises from coordinating efforts across diverse contributors with varying levels of expertise and objectives specific to AI model fine-tuning, validation, and integration. This complexity can lead to issues such as conflicting modifications, redundant work, or difficulties in tracking contributions, which are particularly pertinent in the context of AI model development where nuances in algorithmic implementations and data processing can significantly impact performance and outcomes.
Potential strategies for process management and collaboration could include implementing robust version control systems tailored to AI model development, establishing clear guidelines and best practices for contributors. Additionally, harnessing collaborative tools and platforms tailored specifically to AI model-centric development, equipped with functionalities optimized for tasks such as algorithmic validation, model versioning, and data management, could significantly enhance coordination and communication among contributors.
@@ 6. Conclusion
In the modern digital landscape, open collaborative software development environments have emerged as a significant trend in the field of software engineering. This paper has explored various dimensions of quality assurance within these environments, encompassing process, personnel, and technology aspects. It also analyze the challenges and opportunities with the emerging of LLMs, and looking ahead, there would be more testing techniques empowered by LLMs, as well as testing techniques tailored for LLMs and related scenarios.

### v2

@@ 5.3. Testing and Quality Assurance for Auto-generated Code
With the advancement of LLMs and their remarkable performance in code generation tasks, developers are increasingly relying on these models for various coding and debugging tasks. Recently, there have been reports of groundbreaking developments, such as Microsoft’s creation of an AI programmer named AutoDev (Microsoft, 2024), capable of mastering full-stack skills, which can not only write code and debug, but also train models and even bid for projects on the largest freelancing platform, Upwork.
Automatic code generation is rapidly gaining traction, yet testing practices in this area lag behind. Currently, most approaches utilize code-focused LLMs to assist engineers in writing code, functioning similarly to pair programming in agile methodologies. In this collaborative dynamic, LLMs generate code while engineers review it, relying primarily on traditional testing methods. However, this mismatch between human-driven workflows and AI-driven automation hinders efficiency. Furthermore, challenges such as LLM hallucinations, data poisoning, and other vulnerabilities introduce significant and often hidden risks to code quality.
Functional concerns of auto-generated code. The integration of AI-generated code into open-source projects raises concerns regarding code reliability. While AI models demonstrate impressive capabilities in code generation, there remain uncertainties about the robustness and correctness of the generated code (Ouyang et al., 2023). Liu et al. explores and evaluates the hallucinations in LLM-powered code generation, and categories them into intent conflicting (e.g., local semantic conflicting), context deviation (e.g., generate repetitive statements), and knowledge conflicting (e.g., using un-imported library) (Liu et al., 2024c). It is challenging for LLMs to detect and correct hallucinations through prompting. Therefore, it is crucial to develop specialized techniques for detecting and mitigating these issues. There have been some attempts targeted at specific tasks (Luo et al., 2024), but they have yet to achieve satisfactory results.
Non-functional concerns of auto-generated code. In addition to functional correctness, the non-functional requirements of automatically generated code are also critically important. For example, Zhang et al. proposed EffiBench, a benchmark for assessing the efficiency of automatically generated code (Huang et al., 2024). Similarly, maintainability and security concerns, as highlighted by Asare et al. (Asare et al., 2023), are essential aspects of software quality. Developing robust tools and benchmarks to address these non-functional requirements is crucial for ensuring the reliability and practicality of AI-generated code in real-world applications.
In addition to the inherent weaknesses of LLMs that may result in insecure code, external attackers can also introduce security risks into automatically generated code. For example, model publishers might embed malicious backdoors, causing the model to perform normally on standard inputs but generate harmful outputs (i.e., insecure code) when exposed to specific triggered inputs (Yang et al., 2024). These malicious code could lead to data theft, unauthorized software behavior, or other security breaches. Therefore, testing for automatically generated code must also account for such scenarios to ensure robustness against these threats.
Testing in terms of system-level auto-generated code. The lack of human oversight in the code generation process may lead to the introduction of low-quality code that could compromise the integrity of software systems. The sheer volume of AI-generated contributions necessitates scalable testing processes to ensure compliance with coding standards, adheres to best practices, and complies with project-specific requirements. To meet these needs, it is crucial to develop testing frameworks specifically tailored to address the nuances of auto-generated code. These frameworks should integrate static analysis tools, automated test generation, and dynamic testing techniques to identify and resolve potential issues effectively. Moreover, fostering collaboration between AI developers, software engineers, and open-source maintainers is essential to facilitate the seamless integration of AI-generated contributions while upholding the quality and reliability of open-source software projects.
5.4. Collaborative Testing Between Human and AI
Re-define communication and collaboration between humans and AI systems. As described in Section 3, previous studies on human and AI collaboration mainly focused on using AI-related technologies to provide intelligent services for humans, such as recommending suitable issue reports. However, with the advancement of technologies like LLMs, AI’s capabilities have become more prominent, leading to the emergence of many AI-powered techniques and even AI-powered testers. Therefore, we need to re-define communication and collaboration between humans and AI systems during the software testing and quality assurance process. A widely accepted notion is the human-in-the-loop methodology.
In fact, there are already relevant implementations in the field of software testing. For instance, Zamprogno et al. applied it to test assertion generation, where the developer selects the variables they want assertions for, the tool generates assertions only for these variables, and the developer evaluates the relatively small number of generated assertions, ensuring that only useful assertions are persisted in their test cases (Zamprogno et al., 2022). Similarly, Geethal et al. applied it to program repair, using it in conjunction with active learning techniques to present test cases with a higher probability of being labeled as failing to the human (Geethal et al., 2023). These methods have achieved better performance compared to purely automated techniques.
Meanwhile, the launch of LangGraph, a library to help developers build multi-actor, multi-step, stateful LLM applications, has already supported two “human-in-the-loop” features in OpenGPTs: Interrupt and Authorize (langchain, 2024). The first mode, Interrupt, is the simplest form of control—users monitor the streaming output of the application and manually interrupt it when they deem necessary. The second control mode is Authorize, where users pre-define that the application should hand off control to them whenever a particular actor is about to be called. This underscores the significance of human involvement, and these two modes can inspire the creation of other interaction patterns between humans and AI systems for software testing applications.
@@ 6. Conclusion
The open-collaborative software development paradigm, empowered by internet technology, has significantly transformed the landscape of software testing. This paper explores the interconnected dimensions of process, personnel, and technology in the context of modern software testing. It also examines the challenges and opportunities presented by the emergence of LLMs.
In fact, since the release of ChatGPT on November 30, 2022, the AI field has undergone profound changes. What seemed like forward-thinking ideas just a month ago may now appear commonplace due to the rapid emergence of new technologies. In this era of accelerated productivity, it is difficult to predict what will happen in the next five to ten years. Returning to the field of software engineering and software testing, the development of AI is rapidly transforming related research tasks, and we have witnessed how some research areas that were highly relevant just three years ago have now become less frequently discussed. This roadmap is an attempt to reflect on and analyze the past from the vantage point of the present, offering insights into the challenges and opportunities ahead based on our experience. It is clear that software testing will make significant strides in the future, but no one can predict with certainty what those advancements will look like, which is precisely what makes this era so exciting.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China Grant No.62232016, No.62072442, 62402484, 62402483, Youth Innovation Promotion Association CAS, Basic Research Program of ISCAS Grant No. ISCAS-JCZD-202304, Major Program of ISCAS Grant No. ISCAS-ZD-202302 and ISCAS-ZD-202401, Innovation Team 2024 ISCAS (No. 2024-66).